### v3

@@ Abstract
Understanding people’s preferences is crucial for urban planning, yet current approaches often combine responses from multi-cultural populations, obscuring demographic differences and risking amplifying biases. We conducted a large-scale urban visual perception survey of streetscapes worldwide using street view imagery, examining how demographics – including gender, age, income, education, race and ethnicity, and, for the first time, personality traits – shape perceptions among 1,000 participants with balanced demographics from five countries and 45 nationalities. This dataset, Street Perception Evaluation Considering Socioeconomics (SPECS), reveals demographic- and personality-based differences across six traditional indicators (safe, lively, wealthy, beautiful, boring, depressing) and four new ones (live nearby, walk, cycle, green). Location-based sentiments further shape these preferences. Machine learning models trained on existing global datasets tend to overestimate positive indicators and underestimate negative ones compared to human responses, underscoring the need for local context. Our study aspires to rectify the myopic treatment of street perception, which rarely considers demographics or personality traits.
@@ 1 Introduction
The urban environment fundamentally shapes human perception and behavior in cities [1], with visual elements influencing both immediate perception and deeper cognitive responses to spaces [2]. Urban visual perception – how people interpret and respond to visual information in city environments – has become central to understanding urban experiences [3, 4, 5], particularly as Street View Imagery (SVI) has revolutionized this research domain [6, 7]. Large-scale online surveys using SVI have largely replaced traditional on-site fieldwork [7], significantly expanding geographic scope and participant diversity [6, 8]. Building on this methodological evolution, the MIT Place Pulse projects [9, 6] stand out for their unprecedented scale. Place Pulse 2.0 (PP2) collected ratings for over 110,000 street view images from 56 cities across six continents, evaluated by participants from 162 countries using pairwise comparisons for six perceptual indicators: safe, lively, wealthy, beautiful, boring, and depressing. While machine learning models trained on these comparisons have advanced perception studies in multiple urban regions [5, 10, 11, 12], these multi-city models may introduce bias or mask nuances in subjective preferences.
@@
Our study analyzes urban visual perception across five continents through 1,000 demographically diverse participants from Santiago, Chile; the greater Amsterdam area, Netherlands; Nigeria; Singapore; and San Francisco and Santa Clara, USA, who rated 400 street-level images from their own or neighboring cities, addressing critical gaps in previous research. Unlike most urban perception studies, we prioritized acquiring comprehensive demographic characteristics [29], including gender, age, education, income, race/ethnicity, and, for the first time, personality traits (extraversion, agreeableness, conscientiousness, neuroticism, and openness). Building on previous smaller-scale deployments [30], we expanded both geographic scope and perceptual dimensions by introducing four new indicators (live nearby, walk, cycle, and green) to complement the traditional six (safe, lively, wealthy, beautiful, boring, and depressing).
Our statistical analyses identify how demographic characteristics and personality traits moderate perception across multiple cities, while also examining, for the first time, magnitude bias propagation in widely-used machine learning models trained on the PP2 dataset, which are frequently applied to specific urban contexts such as Shanghai and Beijing [5], Boston and Los Angeles [31], Amsterdam [32], and New York and Singapore [33]. We conducted comparative analyses in two key scenarios: multi-city imagery rated by single-city participants and single-city imagery rated by multi-city participants. Additionally, we quantified how participants rated their own cities compared to others, revealing critical differences in perception when using non-local imagery for localized evaluation. Lastly, we validated our new proposed indicators by analyzing their distinctiveness and complementary relationships with the existing six, highlighting their potential to support ongoing urban research questions on liveability, cyclability, walkability, and greenery.
@@
Figure 1: Research gaps and key questions in urban visual perception studies, followed by our methodology to bridge them. The profile survey included demographics (gender, age group, education level, annual household income, and race and ethnicity) and personality traits (extraversion, agreeableness, conscientiousness, neuroticism, openness). The human perception ratings included the traditional six indicators (safe, wealthy, lively, beautiful, depressing, and boring) and new proposed four indicators (live nearby, walk, cycle, and green) for pairwise comparisons.
@@ 2.1 Demographics and personalities as moderators of perception
Figure 2: Statistical difference of perception scores by demographics. Perception Q scores are calculated from ratings by participants in each group for all (without location grouping) and each location. Locations with no significant differences in any indicator and demographic groups with fewer than $n$ samples (rated images with at least four pairwise comparison by participants per group) are not shown. Significance thresholds *$p<0.05$.
@@
While overall we found fewer differences within demographic interaction nested groups, i.e., a total of four compared to eight in single demographic groups (Figure 2), among all the combinations of ‘gender $\times$ age group $\times$ AHI’ nested groups, half the differences happened within-country and half across all locations (Figure 9). Differences formed by all three demographics combined were found for the lively indicator between “Female x 30-39 x Bottom” and “Male x 21-29 x Middle” nested groups only when all participants’ responses were combined, although under a small sample size ($n\geq 5$, top plot in Figure 9). Within ‘gender $\times$ age group’ nested groups perception differences for the walk indicator between genders of the “Above 50” age group in the USA also under a smaller sample size ($n\geq 5$, second plot from the top Figure 9). Then, differences for the safe indicator between genders of the “Middle” AHI group were found in Netherlands under a larger sample size. Finally, perceptual differences between age groups “21-29” and “30-39” of the “Bottom” AHI bracket were found for the lively indicator across all locations. Both perceptual differences for safe and lively were found with higher sample size ($n\geq 20$, last two bottom plots Figure 9, respectively).
Figure 3: Statistical difference of perception scores by personality traits. Perception Q scores are calculated from ratings by participants in each group for all (without location grouping) and each location. Locations with no significant differences in any indicator and personality groups with fewer than $n$ samples (rated images with at least four pairwise comparison by participants per group) are not shown. Significance threshold *$p<0.05$.
@@ 2.2 Multi-city imagery & single-city participants
Figure 4: Street-level images perception Q scores statistical comparisons based on where people are from (top), where SVIs were taken (middle), and rated by people from their own versus other cities (bottom). Significance threshold *$p<0.05$.
@@
The perception scores given by the ViT-PP2 model do not follow this trend. These scores are overestimated, with median scores 15%, 24%, 38%, and 70% higher than the combined (All) median, in positive indicators – safe, wealthy, wealthy, and beautiful, respectively. Conversely, the scores are underestimated with median scores as low as 32% lower than the combined (All) median, in negative indicators – boring and depressing. The difference in the latter is much lesser, with the combined (All) median Q Score being 1.29% lower than the ViT-PP2 median, and the only difference that is not statistically significant (bottom left subfigure on top plot in Figure 4).
@@ 2.4 Perceptual variation from residents vs non-residents
More than 78% participants in each country have lived in their respective cities – Santiago (Chile), the greater Amsterdam area (Netherlands), Nigeria (multiple cities), Singapore, and San Francisco and Santa Clara (USA) – for more than five years, and less than 5% of them for less than one year (Table 3). This time spent living in the city suggests that the participants would have assimilated into the local culture, customs, and behaviors, and their responses can be regarded as those of someone originally from that area.
@@ 2.5 Perceptual indicators correlation
Figure 5: Correlation between our four new indicators (rows) and the six predominantly used indicators (columns). All correlations are significant ($p<0.05$) except for the correlations of the green indicator and the negative indicators boring and depressing.
We performed a linear correlation analysis and found that our new proposed indicators correctly capture the positive and negative sentiment of the existing perception indicators. The new indicators live nearby, walk, and cycle show a positive correlation (R$>$0.40) with the traditional positive indicators (safe, lively, wealthy, and beautiful) and a negative correlation (R$<$-0.40) with the negative indicators (boring and depressing) (Figure 5). An exception is seen in the relationship between cycle and lively (R$>$0.20) and cycle and boring (R$<$-0.20). The newly proposed green indicator shows an inverse result of negative correlation with a positive indicator (lively) and vice versa (boring). An extended linear correlation analysis is shown in Appendix EFigure 10.
We further investigated non-linear relationships with polynomial regressions. Figure 12 shows the best fitting curve in red (linear, quadratic, or cubic) in terms of $R^{2}$, for the different perceptual pairs, i.e., new indicator vs traditional indicator. In most instances, the best fitting model for live nearby, walk, and cycle is a linear model (in three, five, and four out of the six perceptual pairs, respectively). In the remaining perceptual pairs, the best fitting model has, in all but one case, an $R^{2}$ value 0.1 higher than the respective linear model, with the highest difference being 0.2 in the live nearby vs wealthy pair. This suggests that these three new perceptual indicators are predominately linearly correlated with the traditional six indicators. The green perceptual indicator appears to have a more complex relationship with the traditional six indicators (12(b)). This indicator best correlates linearly with safe and lively, R values of 0.30 and -0.19, respectively and $R^{2}$ values of 0.09 and 0.04. There is a weak non-linear relationship between the green perceptual indicator and wealthy, boring, and depressing indicators ($0.03\leq R^{2}\leq 0.05$), though still three times higher than $R^{2}$ values from linear models. Nevertheless, the relationship between green and beautiful, although relatively low ($R^{2}=0.12$ with a cubic model and $R^{2}$ = 0.11 with a linear model), is as much as three times higher in $R^{2}$ values than the other correlations.
@@ 3 Discussion
Across all cities chosen, the statistical analysis revealed that differences in global urban visual perception are explained by demographics and personalities (Figure 2). While existing work found safety perceptions, in specific locations, are explained by gender [34, 35, 15, 14, 16], our findings revealed a potential aggregated difference when considering all five locations together. Similar to work done in [11], we did not find statistically significant differences between age groups for safety perceptions, although the age group representation in said work was less granular: below or above 50 years old. Another work also did not find significant differences in perception scores by age or gender [9], although similarly to [11], age groups were analyzed as below or above a given age, i.e., 28 years old in said study. Unlike work in [34], we did not find significant differences in the boring indicator for gender, or any, demographic groups. Nevertheless, our findings show that in other indicators, both traditional and new, there are significant differences for all five demographic characteristics in different countries. More importantly, these significant differences are found in cities from continents often overlooked by the research community, e.g., Africa [36]. Furthermore, although nested demographic groups presented limitations in sample size, we found significant differences with reasonable sample sizes ($n\geq 20$) in groups that included AHI (Figure 9).
@@
Our work asserts the importance of multi-city and multi-population analysis. Previous multi-country work either analyzed only a single country of residency of a dataset [9], did not consider participants’ location [6], or focused on a single feature, e.g., biophilia [8]. Equally importantly, our work evidences that targeted intervention should consider the perception of locals in their own city [42], as it will differ greatly if the perceptual evaluation is done using external SVI or by a machine learning model prediction [16]. Nevertheless, we do not claim generalization or universality of these patterns, and we are limited to evaluating a single city in each country as the scope of this work. Our contribution lies in thoroughly, and for the first time, assessing and quantifying these perceptual differences across multiple locations and populations, highlighting the importance of human-centric and context-aware methodologies for future larger urban visual perceptions studies. While we handled inter-country biases by analyzing demeaned perception scores in their respective locations (bottom plot in Figure 4), in-depth analysis on cross-cultural, location-based biases remains a promising research direction. At the same time, intra-country differences should be further analyzed, especially in larger and more diverse countries.
@@
Despite the demographic diversity and geographic location of participants, most of our proposed indicators (live nearby, cycle, and walk) capture, predominantly linearly, the trends of the existing six indicators. These new dimensions are able to expand perception studies with indicators reflecting a user’s willingness to be associated with an urban scene (e.g., live, walk, or cycle in the surroundings). The green indicator introduces the opportunity to expand greenery assessment from traditional objective measurements such as pixel-ratio metrics, e.g., Green View Index (GVI), to subjective context-aware human perception data. While the relationships found with other perceptual indicators were more complex and nonlinear, we found around half of the statistically significant differences in perception scores between demographic and personality profiles (6 out of 13) under these four new indicators. These highlight the potential of these indicators as input for targeted interventions. As different demographic groups in diverse location perceive them differently; e.g., different live nearby and cycle preferences between gender groups in San Francisco and Santa Clara (USA), and Abuja (Nigeria), respectively (Figure 2); different green preferences between high and low agreeableness groups in Singapore; and different walk preferences in genders of a given age group in San Francisco and Santa Clara (USA) (Figure 9). These indicators offer a clearer human subjective perspective that, when complemented with objective measurements, leads to improved urban solutions and assessments [42, 11, 33]. We propose that researchers consider including these indicators where possible and when relevant.
@@
This work contributes to the few studies in multi-city perception, emphasizes the importance of human-centric data collection, and releases a global and diverse perception dataset. Methodologically, we introduced a global and, participants-wise, demographically balanced urban visual perception called SPECS and we thoroughly analyzed perception scores by demographic and personality groups. Moreover, we quantified and statistically verified, for the first time, the magnitude bias of off-the-shelf machine learning models in perceptual predictions despite being trained on a global dataset like PP2, we isolated and analyzed the effect of how people are influenced by preconceived conceptions of urban contexts, and introduced and validated new indicators relevant to the six commonly used. Our results draws us closer to understanding in what contexts, e.g., locations and perceptual indicators, urban visual perception is most affected by demographics and personality. While we are not able to pinpoint a specific number or percentage of how much people’s profiles drive perception responses, we found that as such it can lead to significant differences depending on the profiles and locations. These insights on profiles and location groups can be extended to other urban studies, such as building exterior evaluation for urban development and strategies [43]. Street-level image-based indicators closely correlate with a city’s social structure and economic development [44]. Thus, as SVI captures the human-level perspective, more closely aligned with people’s experiences in cities, studying and understanding urban streetscapes allows researchers and stakeholders to take informed actions [45]. While PP2 remains an impressive effort valuable for urban science, our analysis reveals its limitations in localized contexts, suggesting that generalized or one-size-fits-all models may produce distorted results. We recommend that researchers fine-tune their perception prediction models with local data, even though this may require additional data collection efforts. This approach is likely to produce more accurate and better performing models [11, 16].

### v4

@@ Abstract
Understanding people’s preferences is crucial for urban planning, yet current approaches often combine responses from multi-cultural populations, obscuring demographic differences and risking amplifying biases. We conducted a largescale urban visual perception survey of streetscapes worldwide using street view imagery, examining how demographics – including gender, age, income, education, race and ethnicity, and personality traits – shape perceptions among 1,000 participants with balanced demographics from five countries and 45 nationalities. This dataset, Street Perception Evaluation Considering Socioeconomics (SPECS), reveals demographic- and personality-based differences across six traditional indicators – safe, lively, wealthy, beautiful, boring, depressing – and four new ones – live nearby, walk, cycle, green. Location-based sentiments further shape these preferences. Machine learning models trained on existing global datasets tend to overestimate positive indicators and underestimate negative ones compared to human responses, underscoring the need for local context. Our study aspires to rectify the myopic treatment of street perception, which rarely considers demographics or personality traits.
@@ 1 Introduction
The urban environment fundamentally shapes human perception and behavior in cities [1], with visual elements influencing both immediate perception and deeper cognitive responses to spaces [2]. Urban visual perception – how people interpret and respond to visual information in city environments – has become central to understanding urban experiences [3, 4, 5], particularly as Street View Imagery (SVI) has revolutionized this research domain [6, 7]. Large-scale online surveys using SVI have largely replaced traditional on-site fieldwork [7], significantly expanding geographic scope and participant diversity [6, 8].
Building on this methodological evolution, the MIT Place Pulse projects [9, 6] stand out for their unprecedented scale. Place Pulse 2.0 (PP2) collected ratings for over 110,000 street view images from 56 cities across six continents, evaluated by participants from 162 countries using pairwise comparisons for six perceptual indicators: safe, lively, wealthy, beautiful, boring, and depressing. While machine learning models trained on these comparisons have advanced perception studies in multiple urban regions [5, 10, 11, 12], these multi-city models may introduce bias or mask nuances in subjective preferences.
@@
Our study analyzes urban visual perception across five continents through 1,000 demographically diverse participants from Santiago, Chile; the greater Amsterdam area, Netherlands; Nigeria; Singapore; and San Francisco and Santa Clara, USA, who rated 400 street-level images from their own or neighboring cities, addressing critical gaps in previous research. Unlike most urban perception studies, we prioritized acquiring comprehensive demographic characteristics [29], including gender, age, education, income, race/ethnicity, and personality traits (extraversion, agreeableness, conscientiousness, neuroticism, and openness). Building on previous smaller-scale deployments [30], we expanded both geographic scope and perceptual dimensions by introducing four new indicators (live nearby, walk, cycle, and green) to complement the traditional six (safe, lively, wealthy, beautiful, boring, and depressing).
Our statistical analyses identify how demographic characteristics and personality traits moderate perception across multiple cities, while also examining magnitude bias propagation in widely-used machine learning models trained on the PP2 dataset, which are frequently applied to specific urban contexts such as Shanghai and Beijing [5], Boston and Los Angeles [31], Amsterdam [32], and New York and Singapore [33]. We conducted comparative analyses in two key scenarios: multi-city imagery rated by single-city participants and single-city imagery rated by multi-city participants. Additionally, we quantified how participants rated their own cities compared to others, revealing critical differences in perception when using non-local imagery for localized evaluation. Lastly, we validated our new proposed indicators by analyzing their distinctiveness and complementary relationships with the existing six, highlighting their potential to support ongoing urban research questions on liveability, cyclability, walkability, and greenery.
@@
Figure 1: Research gaps and key questions in urban visual perception studies, followed by our methodology to bridge them. The profile survey included demographics (gender, age group, education level, annual household income, and race and ethnicity) and personality traits (extraversion, agreeableness, conscientiousness, neuroticism, openness). The human perception ratings included the traditional six indicators (safe, wealthy, lively, beautiful, depressing, and boring) and new proposed four indicators (live nearby, walk, cycle, and green) for pairwise comparisons. Source of imagery: Mapillary and KartaView contributors. We acknowledge the use of icons from The Noun Project, created by various authors and licensed under $C$ BY 3.0.
@@ 2.1 Demographics and personalities as moderators of perception
Figure 2: Statistical difference of perception scores by demographics. Perception Q scores are calculated from ratings by participants in each group for all (without location grouping) and each location. Welch’s ANOVA was used for demographic attributes with only two groups and for demographic attributes with more than two groups, we performed the Games-Howell post-hoc test. Minimum sample size $n$ (rated images with at least four pairwise comparison by participants per group) is shown for each demographic profile. Locations with no significant differences in any indicator and demographic groups with fewer than $n$ samples are not shown. Significance thresholds at $p<0.05$.
@@
While overall we found fewer differences within demographic interaction nested groups, i.e., a total of four compared to eight in single demographic groups (Figure 2), among all the combinations of ‘gender $\times$ age group $\times$ AHI’ nested groups, half the differences happened within-country and half across all locations (Extended data Fig. 1). Differences formed by all three demographics combined were found for the lively indicator between “Female x 30-39 x Bottom” and “Male x 21-29 x Middle” nested groups only when all participants’ responses were combined, although under a small sample size ($n\geq 5$, top plot in Extended Data Fig. 1). Within ‘gender $\times$ age group’ nested groups perception differences for the walk indicator between genders of the “Above 50” age group in the USA also under a smaller sample size ($n\geq 5$, second plot from the top Extended data Fig. 1). Then, differences for the safe indicator between genders of the “Middle” AHI group were found in Netherlands under a larger sample size. Finally, perceptual differences between age groups “21-29” and “30-39” of the “Bottom” AHI bracket were found for the lively indicator across all locations. Both perceptual differences for safe and lively were found with higher sample size ($n\geq 20$, last two bottom plots Extended data Fig. 1, respectively).
[…]
@@ 2.2 Multi-city imagery & single-city participants
Figure 4: Street-level images perception Q scores statistical comparisons based on where people are from (top), where SVIs were taken (middle), and rated by people from their own versus other cities (bottom). Welch’s ANOVA was used to compared the combined scores by location (All) against the ML model predictions (top) and to compare scores by each location against the ML model predictions (middle). Minimum sample size $n$ (rated images with at least four (top) or 22 (middle) pairwise comparison by participants per location) is shown. Significance threshold at *$p<0.05$.
@@
The perception scores given by the ViT-PP2 model do not follow this trend. These scores are overestimated, with median scores 15%, 71%, 24%, and 38% higher than the combined (All) median, in positive indicators – safe, lively, wealthy, and beautiful, respectively. Conversely, the scores are underestimated with median scores as low as 32% lower than the combined (All) median, in negative indicators – boring and depressing. The difference in the latter is much lesser, with the combined (All) median Q Score being 1% lower than the ViT-PP2 median, and the only difference that is not statistically significant (bottom left subfigure on top plot in Figure 4).
@@ 2.4 Perceptual variation from residents vs non-residents
More than 78% participants in each country have lived in their respective cities – Santiago (Chile), the greater Amsterdam area (Netherlands), Nigeria (multiple cities), Singapore, and San Francisco and Santa Clara (USA) – for more than five years, and less than 5% of them for less than one year (Supplementary material Table 1). This time spent living in the city suggests that the participants would have assimilated into the local culture, customs, and behaviors, and their responses can be regarded as those of someone originally from that area.
@@ 2.5 Perceptual indicators correlation
Figure 5: Correlation between our four new indicators (rows) and the six predominantly used indicators (columns) across all locations. Pearson correlation R-values where all correlations are significant ($p<0.05$) except for the correlations of the green indicator and the positive indicator wealthy and negative indicators boring and depressing. Minimum sample size $n$ (rated images with at least 22 pairwise comparison by participants per indicator) is shown.
We performed a linear correlation analysis and found that our new proposed indicators correctly capture the positive and negative sentiment of the existing perception indicators. The new indicators live nearby, walk, and cycle show a positive correlation (R>>0.40) with the traditional positive indicators (safe, lively, wealthy, and beautiful) and a negative correlation (R<<-0.40) with the negative indicators (boring and depressing) (Figure 5). An exception is seen in the relationship between cycle and lively (R>>0.20) and cycle and boring (R<<-0.20). The newly proposed green indicator shows an inverse result of negative correlation with a positive indicator (lively) and vice versa (boring). An extended linear correlation analysis is shown in Extended data Fig. 2.
We further investigated non-linear relationships with polynomial regressions. Extended data Fig. 3 shows the best fitting curve in red (linear, quadratic, or cubic) in terms of $R^{2}$, for the different perceptual pairs, i.e., new indicator vs traditional indicator. In most instances, the best fitting model for live nearby, walk, and cycle is a linear model (in three, five, and four out of the six perceptual pairs, respectively). In the remaining perceptual pairs, the best fitting model has, in all but one case, an $R^{2}$ value 0.1 higher than the respective linear model, with the highest difference being 0.2 in the live nearby vs wealthy pair. This suggests that these three new perceptual indicators are predominately linearly correlated with the traditional six indicators. The green perceptual indicator appears to have a more complex relationship with the traditional six indicators (Extended data Fig. 3d). This indicator best correlates linearly with safe and lively, R values of 0.30 and -0.19, respectively and $R^{2}$ values of 0.09 and 0.04. There is a weak non-linear relationship between the green perceptual indicator and wealthy, boring, and depressing indicators ($0.03\leq R^{2}\leq 0.05$), though still three times higher than $R^{2}$ values from linear models. Nevertheless, the relationship between green and beautiful, although relatively low ($R^{2}=0.12$ with a cubic model and $R^{2}$ = 0.11 with a linear model), is as much as three times higher in $R^{2}$ values than the other correlations.
@@ 3 Discussion
Across all cities chosen, the statistical analysis revealed that differences in global urban visual perception are explained by demographics and personalities (Figure 2). While existing work found safety perceptions, in specific locations, are explained by gender [34, 35, 15, 14, 16], our findings revealed a potential aggregated difference when considering all five locations together. Similar to work done in [11], we did not find statistically significant differences between age groups for safety perceptions, although the age group representation in said work was less granular: below or above 50 years old. Another work also did not find significant differences in perception scores by age or gender [9], although similarly to [11], age groups were analyzed as below or above a given age, i.e., 28 years old in said study. Unlike work in [34], we did not find significant differences in the boring indicator for gender, or any, demographic groups. Nevertheless, our findings show that in other indicators, both traditional and new, there are significant differences for all five demographic characteristics in different countries. More importantly, these significant differences are found in cities from continents often overlooked by the research community, e.g., Africa [36]. Furthermore, although nested demographic groups presented limitations in sample size, we found significant differences with reasonable sample sizes ($n\geq 20$) in groups that included AHI (Extended data Fig. 1).
@@
Our work asserts the importance of multi-city and multi-population analysis. Previous multi-country work either analyzed only a single country of residency of a dataset [9], did not consider participants’ location [6], or focused on a single feature, e.g., biophilia [8]. Equally importantly, our work evidences that targeted intervention should consider the perception of locals in their own city [42], as it will differ greatly if the perceptual evaluation is done using external SVI or by a machine learning model prediction [16]. Nevertheless, we do not claim generalization or universality of these patterns, and we are limited to evaluating a single city in each country as the scope of this work. Our contribution lies in thoroughly assessing and quantifying these perceptual differences across multiple locations and populations, highlighting the importance of human-centric and context-aware methodologies for future larger urban visual perceptions studies. While we handled inter-country biases by analyzing demeaned perception scores in their respective locations (bottom plot in Figure 4), in-depth analysis on cross-cultural, location-based biases remains a promising research direction. At the same time, intra-country differences should be further analyzed, especially in larger and more diverse countries.
@@
Despite the demographic diversity and geographic location of participants, most of our proposed indicators (live nearby, cycle, and walk) capture, predominantly linearly, the trends of the existing six indicators. These new dimensions are able to expand perception studies with indicators reflecting a user’s willingness to be associated with an urban scene (e.g., live, walk, or cycle in the surroundings). The green indicator introduces the opportunity to expand greenery assessment from traditional objective measurements such as pixel-ratio metrics, e.g., Green View Index (GVI), to subjective context-aware human perception data. While the relationships found with other perceptual indicators were more complex and nonlinear, we found around half of the statistically significant differences in perception scores between demographic and personality profiles (6 out of 13) under these four new indicators. These highlight the potential of these indicators as input for targeted interventions. As different demographic groups in diverse location perceive them differently; e.g., different live nearby and cycle preferences between gender groups in San Francisco and Santa Clara (USA), and Abuja (Nigeria), respectively (Figure 2); different green preferences between high and low agreeableness groups in Singapore; and different walk preferences in genders of a given age group in San Francisco and Santa Clara (USA) (Extended data Fig. 1). These indicators offer a clearer human subjective perspective that, when complemented with objective measurements, leads to improved urban solutions and assessments [42, 11, 33]. We propose that researchers consider including these indicators where possible and when relevant.
@@
This work contributes to the few studies in multi-city perception, emphasizes the importance of human-centric data collection, and releases a global and diverse perception dataset. Methodologically, we introduced a global and, participants-wise, demographically balanced urban visual perception dataset called Street Perception Evaluation Considering Socioeconomics (SPECS) and we thoroughly analyzed perception scores by demographic and personality groups. Moreover, we quantified and statistically verified the magnitude bias of off-the-shelf machine learning models in perceptual predictions despite being trained on a global dataset like PP2, we isolated and analyzed the effect of how people are influenced by preconceived conceptions of urban contexts, and introduced and validated new indicators relevant to the six commonly used. Our results draws us closer to understanding in what contexts, e.g., locations and perceptual indicators, urban visual perception is most affected by demographics and personality. While we are not able to pinpoint a specific number or percentage of how much people’s profiles drive perception responses, we found that as such it can lead to significant differences depending on the profiles and locations. These insights on profiles and location groups can be extended to other urban studies, such as building exterior evaluation for urban development and strategies [43]. Street-level image-based indicators closely correlate with a city’s social structure and economic development [44]. Thus, as SVI captures the human-level perspective, more closely aligned with people’s experiences in cities, studying and understanding urban streetscapes allows researchers and stakeholders to take informed actions [45]. While PP2 remains an impressive effort valuable for urban science, our analysis reveals its limitations in localized contexts, suggesting that generalized or one-size-fits-all models may produce distorted results. We recommend that researchers fine-tune their perception prediction models with local data, even though this may require additional data collection efforts. This approach is likely to produce more accurate and better performing models [11, 16].